OpenStudy (anonymous):

What is the standard form of the equation of a circle with its center at (2, -3) and passing through the point (-2, 0)?

Nnesha (nnesha):

standard form equation of circle is \[\huge\rm (x-h)^2+(y-k)^2=r^2\] where (h,k) is the center and substitute (x,y) for (-2,10) to find radius

Nnesha (nnesha):

let me know if you din't get it

Nnesha (nnesha):

didn't*

OpenStudy (anonymous):

Ya I don't sorry

Nnesha (nnesha):

alright so center point is (2,-3) which is (h,k) and (x,y) point is (-2,0) so replace (h,k) and x y with their values

Nnesha (nnesha):

\[\huge\rm (\color{ReD}{h},\color{blue}{k}),(\color{reD}{2},\color{blue}{-3})\] first coordinate is h 2nd one is k

Nnesha (nnesha):

\[\huge\rm (x-h)^2+(y-k)^2=r^2\] so replace h with 2 and k with -3

Nnesha (nnesha):

(-2, 0) what is x and y in this order pair ?

OpenStudy (anonymous):

-3,2

Nnesha (nnesha):

that's the center point in this order piar (-2,0) which one is x-coordinate or which one is y-coordinate

Nnesha (nnesha):

?

OpenStudy (anonymous):

2 is the x

Nnesha (nnesha):

it's -2 and yes now substitute \[\huge\rm (-2-2)^2+(0-(-3))^2=r^2\] solve for r

OpenStudy (anonymous):

-7?

Nnesha (nnesha):

nope i didn't get that

Nnesha (nnesha):

solve for r^2 **